ISP-Hookup-HOWTO

Egil Kvaleberg, egilk@sn.no v1.23, 9 December 1996 This document describes how to use Linux to connect to an Internet Service Provider via a dial-up modem connection. As well as the basic dial-up procedure and IP establishment, email and news handling is covered.

Contents

1 Introduction 2 1.1 New versions of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Feedback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Disclaimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Copyright information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 How do I connect to the rest of the world? 4 2.1 The basic configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 How do I surf ? 6

4 How do I send and receive email? 7

5 News 8 5.1 How do I set up an online news-reader? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5.2 How do I set up an offline news-reader? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.3 How do I set up C News? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.4 How do I set up Leafnode? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6 How do I automate the connection procedure? 12

7 Final words 15 7.1 Other things I should know about? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

8 ISP specific information 16 8.1 How do I learn more? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 8.2 Thanks to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

1 Introduction

This description has been made to answer a few questions about how dial-up ISP (Internet Service Provider) subscribers may configure and use Linux. To aid those who will connect their Linux based machines to an ISP for the first time, an attempt has been made to cover most issues encountered. This quite unavoidably will create a certain degree of overlap with other Linux Howto-documents and LDP books. Reference should be made to these documents to provide better understanding and detail. Much of the existing documentation is targeted towards users with a certain degree of experience, and first time users will often have trouble sorting out the relevant information. To simplify, the examples given will assume the following:

• User name: dirk

   • Password: PrettySecret

• Internet service provider: acme.net

   • Email server: mail.acme.net
   • News server: news.acme.net

• Name server: 193.212.1.0

• Phone number: 12345678

Our dirk will be calling his machine roderick. All references in the table above should naturally be replaced by whatever is valid for the ISP one is using. Often, just a minimum of changes will otherwise be required for users with different ISPs. I would like to be informed about what problems you encounter on this account.

1.1 New versions of this document

New versions of this document will be periodically posted to comp.os.linux.answers. They will also be added to the various anonymous FTP sites who archive such information, including:

ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO

In addition, you should generally be able to find this document on the Linux WorldWideWeb home page via:

http://sunsite.unc.edu/mdw/linux.html

Finally, the very latest version of this document should also be available in various formats from:

ftp://ftp.sn.no/user/egilk/ISP-Hookup-HOWTO.txt
ftp://ftp.sn.no/user/egilk/ISP-Hookup-HOWTO.ps.gz

ftp://ftp.sn.no/user/egilk/ISP-Hookup-HOWTO.tar.gz
http://home.sn.no/ egilk/ISP-Hookup-HOWTO.html

1.2 Feedback

All comments, error reports, additional information and criticism of all sorts should be directed to:

egilk@sn.no
http://home.sn.no/ egilk/

1.3 Disclaimer

No liability for the contents of this documents can be accepted. Use the concepts, examples and other content at your own risk. Additionally, this is an early version, with many possibilities for inaccuracies and errors.

One of many possible setups will be described. In the Linux world, there is usually a number of ways in which to accomplish things. Paragraphs containing hints to alternatives are marked by ALT: Please also note that FTP-references often will change slightly as new versions of programs arrive.

As far as I know, only programs that under certain terms may be used or evaluated for personal purposes will be described. Most of the programs will be available complete with source under GNU-like terms.

1.4 Copyright information

This document is copyrighted (c)1996 Egil Kvaleberg and distributed under the following terms:

• Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

• All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

• If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at

gregh@sunsite.unc.edu Finger for phone number and snail mail address.

2 How do I connect to the rest of the world?

It will be assumed that we have installed the essential networking software modules (e.g. essential parts of the Slackware N-series), and that you have set up which serial port that is to be used for the /dev/modem.

The default configuration will usually only allow direct access to /dev/modem as user root.

To connect to ISP shell accounts directly, and to experiment with connection sequences, you may use the minicom program. It is pretty straight forward to use.

2.1 The basic configuration

Configuration of the machine for use on the net should be done as user root. Before proceeding any further, ensure that the file /etc/hosts.deny contains the following line:

     ALL: ALL

You would normally want to allow yourself, so add the following line to /etc/hosts.allow:

     ALL: LOCAL

Or if you insist:

     ALL: 127.0.0.1

For the following, note that it is meant for those connected via PPP and with a dynamic IP address. If you have the benefit of a fixed connection, there will be some differences.

It is nice to have a name connected to the machine, a name that the dynamic IP user really can select as he or she pleases. Put the name in /etc/HOSTNAME:

     roderick

The next step is to set up the name server in /etc/resolv.conf:

     search .
     nameserver 193.212.1.0

The name server must be specified by a numeric IP address, and will be different from ISP to ISP. If required, you can have up to three different servers. They will be requested in the sequence in which they are listed.

If you want to be able to use names like somemachine as an abbreviation for somemachine.acme.net, you must replace the first line with:

     search acme.net

A certain minimum of configuration will also be required in /etc/hosts. Most users will be able to manage with:

     127.0.0.1 localhost
     0.0.0.0   roderick

Those with a fixed IP-address will obviously replace 0.0.0.0 with this.

Likewise, a minimum /etc/networks is:

     loopback  127.0.0.0
     localnet  0.0.0.0

You should also set your external mail domain in /etc/mailname:

     acme.net

The username and password at the ISP must be specified in /etc/ppp/pap-secrets

     dirk * PrettySecret

For those ISPs using CHAP instead of PAP, the filename is /etc/ppp/chap-secrets.

Finally, the nitty gritty regarding the connection procedure itself must be specified before PPP can be initiated. This is done in /etc/ppp/chatscript:

     TIMEOUT 5
     "" ATZ
     OK ATDT12345678
     ABORT "NO CARRIER"
     ABORT BUSY
     ABORT "NO DIALTONE"
     ABORT WAITING
     TIMEOUT 45
     CONNECT ""
     TIMEOUT 5
     "name:" ppp

Details here may have to be tuned somewhat. The phone number in the third line must of course be set as required. Some users may need to replace the ATZ modem initialization string with something more tailored for the modem being used. The last line specifies that one is expecting the prompt name:, and that the response should be ppp when it arrives. Other systems may have other login procedures.

To actually initiate a call, the PPP-protocol may be initiated by issuing the following command:

     exec pppd connect \
          'chat -v -f /etc/ppp/chatscript' \
           -detach crtscts modem defaultroute \
           user dirk \
           /dev/modem 38400

We should now be on-air, and stay up until the program is killed by typing a Ctrl-C. Any messages concerning the connection will be appended to the system logs. To read them, try:

     tail /var/adm/messages

As long as PPP is up, you will have direct access to the Internet, and may use programs like ftp, ncftp, rlogin, telnet, finger etc. All these programs should be part of the network package. Further information concerning PPP is also available from:

/usr/lib/ppp/README.linux
/usr/lib/ppp/README.linux-chat

Finally, an additional word about safety. The file /etc/inetd.conf lists all services that your machine will offer externally. With the /etc/hosts.deny file we have made, no external access will be allowed. For those who need it, access must be allowed explicitly in /etc/hosts.allow. Local traffic may be allowed by:

     ALL: LOCAL
See also man 5 hosts access.

A final small issue: A certain confusion exists regarding the names of the POP-protocols. A definition in /etc/services compatible with just about everything is:

     pop2      109/tcp   pop-2          # PostOffice V.2
     pop3      110/tcp   pop-3 pop # PostOffice V.3

ALT: Instead of chatscript, one might use the much more flexible dip. But not in connection with diald. ALT: Those fortunate enough to have a permanent TCP/IP connection via e.g. an Ethernet may safely ignore anything about PPP, and rather start concentrating about setting up their network card. ALT: Others may not have the possibility of using PPP, but may be able to use SLIP instead, for which there is support in much the same manner as for PPP. Another possibility is UUCP. Others again may have to rely on exchange of news and email be means of SOUP. A description for the latter case may be found in:

ftp://ftp.sn.no/user/bjorn/Linux-offline.tgz

The TERM program is also an option. Refer to the Term-HOWTO.

3 How do I surf ?

If you think that text is the most important, you might want to use the Lynx Web-browser. It is available from:

ftp://sunsite.unc.edu/pub/Linux/system/Network/info-systems/lynx-2.3.bin2.tar.gz

If you have installed X-windows, you can also use one of the many graphical browsers. Chimera may be found at:

ftp://sunsite.unc.edu/pub/Linux/system/Network/info-systems/chimera-1.65.bin.ELF.tar.gz
http://www.unlv.edu/chimera/

Mosaic:

ftp://sunsite.unc.edu/pub/Linux/system/Network/info-systems/Mosaic-2.7b1-aout.tgz
ftp://ftp.NCSA.uiuc.edu/Web/Mosaic/Unix/binaries/2.6

Mozilla (Netscape):

ftp://sunsite.unc.edu/pub/Linux/system/Network/info-systems/netscape-v11b3.tar.gz
ftp://ftp.cs.uit.no/pub/www/netscape

These browsers are constantly available in new and in various ways exciting versions. Use and evaluation of these programs is subject to certain terms. Please observe them.

4 How do I send and receive email?

First of all, ensure that sendmail is installed. Sendmail sorts internal and out-bound mail, and will buffer out-bound mail until such time it is possible to forward it. Sendmail is based on a configuration found in /etc/sendmail.cf. An example suitable for ISP users can be found in: ftp://ftp.sn.no/user/egilk/sendmail.cf This is based on procmail as a delivery agent, but may easily be changed to use deliver. It is if course required to have an official domain address for out-bound mail, something which is specified

in /etc/sendmail.cf:
     # who I masquerade as (null for no masquerading)
     DMacme.net

This assumes that you have the same user name locally as you have at your ISP. Sendmail is now configured for sending directly to the recipient. To avoid long and repeated connections in those cases where the connection to the receiving end is slow and irregular, is is usually nice to use ones ISP as a buffer store. This can be specified by the DS specification:

     # "Smart" relay host (may be null)
     DSmail.acme.net

Beware that sendmail is somewhat sensitive to handling of tab stop characters in sendmail.cf. You might want to use the vi editor to ensure that these tab characters are retained unchanged. Email reception can often be performed via the POP3 protocol, which can be initiated every time the connection is brought up. A script for testing this is:

     sendmail -q
     popclient -3 -v mail.acme.net -u dirk -p "PrettySecret" \
            -k -o /usr/spool/mail/dirk

This script may be started after PPP connection has been established. Beware that this script is just for testing, so ensure that the local mailbox is left untouched while it runs. The -k option means that the mail is kept in the ISP mailbox, and you are simply given a copy of the mail. You would of course want to remove this option once you are confident that your setup is working. Beware that the password will show on the command line. This really should be fixed ASAP. A safe and better version of this script may be found at:

ftp://ftp.sn.no/user/egilk/pop-script.tar.gz

This version of the script requires that procmail is installed, but that is something you’ll never regret anyway:

ftp://sunsite.unc.edu/pub/Linux/system/Mail/mailhandlers/procmail-3.10-2.tar.gz

Procmail is a simple and flexible tool that can sort incoming email based on a large range of criteria. In addition to being able to handle automated tasks like vacation messages and such. Note that when we use procmail directly as in this case, the situation is somewhat different from what is described in the procmail documentation. A .forward is not required, and we also don’t need a .procmailrc. The latter is only required if we want to sort the mail. The user interface for reading and sending of email can be found in programs like Pine or Elm. ALT: For an ordinary dial-up ISP user it is not really necessary to have the sendmail daemon active. To reduce the resource usage, one may thus comment out any startup of sendmail, as is usually found in

/etc/rc.d/rc.M.

ALT: In place of sendmail one might use the simpler smail. You’ll find a good description of it (as well as most other things mentioned here) in the Linux Network Administrator’s Guide. ALT: There is also an m4 macro package for making a fresh sendmail.cf. For a simple installation it might be just as well to modify an existing configuration. ALT: There are also simpler although less flexible alternatives. Pine may run stand-alone as long as it is configured properly, for instance. It might even be possible to use newer versions of some web-browsers. ALT: Many are very enthusiastic regarding the Emacs companion Gnus as an email and news handler. Further information can be found at:

http://www.ifi.uio.no/ larsi/

ALT: An alternative to popclient is pop-perl5. It is available from:

ftp://sunsite.unc.edu/pub/Linux/System/Mail/pop-perl5-1.1.tar.gz

5 News

5.1 How do I set up an online news-reader?

As long as PPP is active, it will be possible to read news online. There are lots of available programs. Two simple alternatives are rtin and trn. To start reading news, the only thing required in terms of configuration in most cases is to do (usually once and for all in the file .profile):

     export NNTPSERVER=news.acme.net

To get the From-address correct in postings, some programs may require:

     export NNTP_INEWS_DOMAIN=acme.net

5.2 How do I set up an offline news-reader?

To be able to read news while offline, and thus reduce phone bills and give greater flexibility, one must set up a local news-spool of one sort or the other. This requires some configuration, and there will also be a certain amount of disk space involved. After initial setup, things should run more or less by themselves, with only some attention needed from time to time.

Two different solutions will be described here.

5.3 How do I set up C News?

The solution described here is based on the news-server C News and the NNTP protocol. C News was originally targeted towards another sort of configuration, but is flexible enough to handle our situation too. One might also use the more recent INN news server, but it might require a bit more in terms of resources. Any way, be careful not to install both; they don’t live together easily.

It is crucial that all maintenance of news is done while logged in as user news, and that all configuration files is placed in /usr/lib/news. One way of handling this is, while logged in as root to write su news;

cd.

The most important files in the configuration are:

• active is an overview over active newsgroups. It is updated as required by the command addgroup,

     e.g. addgroup comp.os.linux.networking y.

• organization should simply contain whatever you want in the Organization: header field, in our case:

     Dirk Gently's Holistic Detective Agency

• mailname should in our case be set to acme.net.

• whoami is set to the name of your site in the Path: thread. In a setup as described here, using NewsX, this name will never leave the machine, so you can set this to whatever you like as long as you are pretty sure it is unique. In this case roderick.

• sys controls fetching and further distribution of news. We will assume the ISP in our case adds acme.net to the Path, and that this is the only news source we have. The example given really tells that we will accept everything that arrives, and that we will only post news to acme.net that it hasn’t seen before, and is originally posted at our own site. In this simplified setup we assume that the all groups will come from a single source. /all specifies the distribution, and must be included. The letter F says that (pointers to) outgoing news articles will be collected in a file.

     ME:all/all::
     acme/acme.net:all,!junk/all:FL:

• A subdirectory for the outgoing news must be created, in our case:

     mkdir /var/spool/news/out.going/acme

• mailpaths controls posting in moderated groups, although this task may usually be left to the ISP.

C News needs a certain degree of daily maintenance, but this can be specified once and for all via the command crontab -e issued as user news. A suggested setup follows; it can be tuned as required:

     # maintain incoming and outgoing batches
     10,40 *  * * * /usr/lib/newsbin/input/newsrun
     # expire C News, once a day
     30 0  * * * /usr/lib/newsbin/expire/doexpire
     # monitor and report if needed
     00 2  * * sat /usr/lib/newsbin/maint/addmissing
     40 3  * * * /usr/lib/newsbin/maint/newswatch
     50 3  * * * /usr/lib/newsbin/maint/newsdaily

newsrun moves articles in and out (twice every hour), doexpire will delete articles as they get old (every night at 00:30), and the three last commands does various supervisory and error correcting tasks.

One should also ensure that things are cleaned up when starting the machine. As user root, add the

following line to /etc/rc.d/rc.local:
     su news -c /usr/lib/newsbin/maint/newsboot

News may be collected via the program NewsX, picking news from an NNTP-server. The program can be found at:

ftp://sunsite.unc.edu/pub/Linux/system/News/newsx-0.1.tar.gz Or:
ftp://ftp.sn.no/user/egilk/newsx-0.3.tar.gz

Setting up NewsX is quite simple. Installation is a classic case of:

     make
     su
     make install
     exit

With the setup outlined here, all you have to do is to create the groups you want to read using the ”addgroup” command.

To fetch articles, user news issues the following commands (assuming communication via PPP or similar is up):

     newsrun
     newsx acme news.acme.net
     newsrun

The option -d gives continuous printout to the screen. Refer to the NewsX documentation for further information.

NewsX will also take care of posting of outgoing news.

To control disposal of articles as they get old, a file explist is required. The comments in this example should explain what we want to do:

     # hold onto history lines 14 days, nobody gets {$>$}120 days
     /expired/           x    14   -
     /bounds/            x    0-1-120 -
     # retain these for 2 months
     comp.sources,comp.os.linux.all     x    60   -
     # noise gets thrown away fast
     junk,control             x    2    -
     # default:  14 days, no archive
     all                 x    14   -

ALT: In a small news-spool, one will often not need the newsgroup control. The traffic is huge compared to the usefulness. The main point is that articles will be canceled, and that groups may be created automatically. To ensure that control messages containing newgroup not shall mess up things for us, a file called newgroupperm specifies what we will allow:

     comp.os.linux  tale@uunet.com yv
     all       any       nq

In this example, all proper groups under comp.os.linux will be created (y), and the user news will be notified (v). Everything else will be silently (q) ignored (n). The last line is sufficient if you want to create all groups manually.

ALT: Alternatives to NewsX are suck, or slurp combined with postit. Slurp uses the NNTP NEWNEWS which will put severe loads on many news server.

5.4 How do I set up Leafnode?

A different solution altogether is to install the integrated package leafnode. This will handle all tasks required for a personal news spool, and is easy to configure. It is available via:

http://www.troll.no/freebies/leafnode.html

As for C News, all news maintenance should be performed as user news.

The home directory for leafnode is in /usr/lib/leafnode. To install, write:

     cd /usr/lib/leafnode
     tar -xzvf leafnode-0.8.tgz
     cd leafnode-0.8
     make
     su
     make install

While still being logged in as root, change the line that controls NNTP in /etc/inetd.conf:

     nntp  stream  tcp  nowait  news  /usr/sbin/tcpd /usr/local/sbin/leafnode

Activate it by:

     killall -HUP inetd

Return to user news by writing exit. In /usr/lib/leafnode/config change the line that defines the NNTP server. In our case:

     server = news.acme.net

Leafnode will look after itself by adding the following command via crontab -e as user news:

     # expire Leafnode, once a day
     0 4 * * * /usr/local/sbin/texpire

The news exchange is also done as user news by the following command (assuming PPP is up and running):

     /usr/local/sbin/fetch

Users who wants to read news should then use the recipe in How do I set up an online news-reader?, except that they configure for the local machine, i.e:

     export NNTPSERVER=localhost

That should be all there is to it. The first fetch will transfer a list of available newsgroups. Leafnode will then monitor what groups the users are requesting, and adapt to this the next time it is activated. ALT: An alternative to leafnode is nntpcache, available from:

ftp://ftp.suburbia.net/pub/nntpcache/nntpcache.tgz

6 How do I automate the connection procedure?

Automated handling of news and email is quite easy to implement in Linux. First and foremost one should make a /usr/lib/ppp/ppp-on that initiates the ISP connection. Often, this file will simply contain the following:

     /usr/sbin/pppd

Further specification will be performed in /etc/ppp/options:

     connect "/usr/lib/ppp/chat -v -f /etc/ppp/chatscript"
     crtscts
     modem
     defaultroute
     asyncmap 00000000
     user dirk
     /dev/modem 38400

To end a connection, use the supplied version of /usr/lib/ppp/ppp-off.

Having tested the functionality of these two scripts, one must then write scripts that perform the various tasks. The script to collect email has been described before, and we will here assume it is located at

/home/dirk/pop.

A script for exchange of email can then be produced in /root/mail:

     #! /bin/sh
     #
     # exchange mail
     # 10 minutes timeout:
     TIMEOUT=600
     DT=10
     # kick sendmail:
     sendmail -q \&
     # retrieve mail:
     su dirk -c /home/dirk/pop
     # wait for sendmail to terminate:
     t=0
     while ! mailq | grep -q "Mail queue is empty"; do
         t=$[$t+$DT]
         if [ $t -gt $TIMEOUT ] ; then
          echo "sendmail -q timeout ($TIMEOUT).."
          exit 1
         fi
         sleep $DT
     done
     exit 0

The script to exchange news may be placed in /usr/lib/news/news:

     #!/bin/sh

6. How do I automate the connection procedure? 14

#
# exchange news
# must be run as news:
cd /usr/lib/news
#update the outgoing batch (C News):
/usr/lib/newsbin/input/newsrun {$<$} /dev/null
#exchange news:
/usr/lib/newsbin/newsx acme news.acme.net
#and flush the incoming batch:
/usr/lib/newsbin/input/newsrun {$<$} /dev/null

A script to connect the various bits and pieces remains, and can be placed in /root/news+mail:

#!/bin/sh
#
# exchange news and email
# must be run as root
#
if ! /usr/lib/ppp/ppp-on; then
    exit 1
fi
trap "/usr/lib/ppp/ppp-off" 1 2 3 15
#exchange news+mail:
/root/mail \&
su news -c ~news/news
wait
#disconnect..
/usr/lib/ppp/ppp-off
#update the incoming batch (C News):
su news -c /usr/lib/newsbin/input/newsrun {$<$} /dev/null \&
exit 0

It is quite easy to make an extension to the above that only will establish a connection if outgoing email and news is present. Lets call it /root/news+mail.cond, and keep in mind that the name of the outgoing news-spool must be updated to suit:

#!/bin/sh
#
# exchange news and email, only if outgoing news or mail
# (C News spool)
if [ -s /var/spool/news/out.going/acme/togo ] ||

         ! ( mailq | grep -q "Mail queue is empty"); then
          /root/news+mail
     fi

The only thing remaining is to specify when all this is going to happen. This is done using the command crontab -e as root. Let us assume that we always want to exchange news and mail at 07:00 in the morning, and after that every 4th hour assuming there are outgoing email and news:

     00 7      * * *     /root/news+mail
     00 11,15,19,23 * * *     /root/news+mail.cond

Ensure that every component is tested well before you connect them together. One may later add several other tasks, such as adjustment of the time of day (using ntpdate), and automatic update (mirroring) of locally maintained WWW and FTP files up to the ISP (using make and ftp).

ALT: Depending on ones preferences, it is also possible to turn the process upside down. Every time a PPP link is initiated, the script /etc/ppp/ip-up will be started. One may here add whatever magic is required to start exchange of email and news. See man pppd for further detail.

ALT: It is also possible to automatically connect PPP whenever network traffic is detected. This is in many ways the more elegant solution, but it is quite dependent on a good configuration to avoid frequent (and costly) connections being made. More information can be found at:

http://www.cs.toronto.edu/ schenk/diald.html

The diald utility is available from:

ftp://sunsite.unc.edu/pub/Linux/system/Network/serial/diald-0.13.tar.gz

At the same location one will also find other variations on the theme PPP connections.

7 Final words

7.1 Other things I should know about?

• Various error messages in the system will normally be issued as internal email. To ensure that these will actually be read, one should create an /etc/aliases. Remember the command newaliases every time you change this. An example that should cover most eventualities is:

     PostMaster: root
     ftp: root
     news: root
     usenet: root
     FaxMaster: root
     fax: root
     WebMaster: root
     MAILER.DAEMON: root

• Many programs for Linux may be found at Sunsite, which is usually quite busy. But there are many mirrors, and every time there is a reference to ftp://sunsite.unc.edu/pub/Linux/.. one should

     try to use a mirror close to home, e.g. ftp://ftp.nvg.unit.no/pub/linux/sunsite/...

• If you happen to be migrating from Yarn, it should be possible to convert these to standard folders using the yarn2mf available at:

ftp://ftp.sn.no/user/egilk/yarn2mf.zip

8 ISP specific information

More specific information for certain ISPs is available from a variety of sources: Demon Internet (demon.co.uk)

ftp://ftp.demon.co.uk/pub/unix/linux/Demon/slack3.0.help.tgz

Easynet TBA Netcom http://www.netcom.com/bin/webtech/NetCruiser/Operating Systems/Linux/linux.cfg.html PowerTech, Schibstednett, Telenor Online

http://home.sn.no/ egilk/no-isp.html

Primenet TBA Stanford

http://www-leland.stanford.edu/ wkn/Linux/network/network.html

If you know of ISP specific information not listed here, please get in touch.

8.1 How do I learn more?

The Linux Documentation Project book called Linux Network Administrator’s Guide by Olaf Kirch is pretty mandatory for anyone that will set up and maintain anything involving TCP/IP and Internet:

ftp://sunsite.unc.edu/pub/Linux/docs/linux-doc-project/network-guide/nag-1.0.ascii.tar.gz

The documentation that follows each software package will normally give you all the detailed information you need, if not always the overview. The man-pages will be the first place to look. Try for instance:

     man pppd

You will also find some documentation about certain programs in the /usr/doc tree, although this is not always well structured. The following HOWTOs will be highly relevant:

• Installation-HOWTO will get the basics sorted.

8. ISP specific information 17

• NET-2-HOWTO is a very thorough description of installation and setup of the NET code. Much of this should already have been done if you use a standard Linux distribution (e.g. Slackware, RedHat, Debian). But many sections on setup and troubleshooting will be very worthwhile.

• Mail-HOWTO explains how to configure various tools. Again, much of this will already have been done for you when you install a standard Linux distribution.

• News-HOWTO is for setting up a (conventional) news spool.

• Tiny-News covers yet another alternative for collecting news.

• PPP-HOWTO is a good description of problems you may encounter when setting up a PPP connection.

• Serial-HOWTO contains everything you need to know about setting up serial ports.

• Mail-Queue tells you how to send up sendmail to always queue remote mail but deliver local mail at once.

8.2 Thanks to

Information here is collected from many sources. Thanks to the following that either indirectly or directly have contributed:

Adam Holt {$<$}holt@graphics.lcs.mit.edu{$>$}
Arne Coucheron {$<$}arneco@oslonett.no{$>$}
Arne Riiber {$<$}riiber@oslonett.no{$>$}
Arnt Gulbrandsen {$<$}agulbra@troll.no{$>$}
Bjorn Steensrud {$<$}bjornst@powertech.no{$>$}
Gisle Hannemyr {$<$}gisle@a.sn.no{$>$}
Hans Amund Rosbach {$<$}haro@sesam.dnv.no{$>$}
Hans Peter Verne {$<$}hpv@ulrik.uio.no{$>$}
Harald T Alvestrand {$<$}Harald.T.Alvestrand@uninett.no{$>$}
Harald Terkelsen  {$<$}Harald.Terkelsen@adm.hioslo.no{$>$}
Haavard Engum {$<$}hobbes@interlink.no{$>$}
Johan S. Seland {$<$}johanss@sn.no{$>$}
John Phillips {$<$}john@linux.demon.co.uk{$>$}
Jorn Lokoy {$<$}jorn@oslonett.no{$>$}
Kenneth Tjostheim {$<$}kenneth.tjostheim@asplanviak.no{$>$}
Kjell M. Myksvoll {$<$}kjell.myksvoll@fou.telenor.no{$>$}
Kjetil T. Homme {$<$}kjetilho@math.uio.no{$>$}
Michael Meissner {$<$}meissner@cygnus.com{$>$}
N J Bailey {$<$}N.J.Bailey@leeds.ac.uk{$>$}
Nicolai Langfeldt {$<$}janl@math.uio.no{$>$}
Ove Ruben R Olsen {$<$}Ove.R.Olsen@ub.uib.no{$>$}
R. Bardarson {$<$}ronb@powernet.net{$>$}
Steinar Fremme {$<$}steinar@fremme.no{$>$}
Sverre H. Huseby {$<$}sverrehu@ifi.uio.no{$>$}
Trond Eivind Glomsrod {$<$}teg@stud.imf.unit.no{$>$}

8. ISP specific information 18

Tommy Larsen {$<$}tommy@mix.hive.no{$>$}